# ReliableMath: Benchmark of Reliable Mathematical Reasoning on Large Language Models

### v1

@@ 3.3 Unsolvable Data Construction Workflow
To obtain high-quality unsolvable problems using the rewriting methods in Sec. 3.2, further professional verification and refinement are required. Prior work on constructing unsolvable and unknown problems in knowledge-intensive tasks has relied exclusively on manual collection or filtering (amayuelas-etal-2024-knowledge, ; yin_large_2023, ). However, mathematical problems present greater challenges, and fully manual rewriting and validation are prohibitively expensive. Therefore, we design a three-stage data construction workflow to ensure both efficiency and quality, as illustrated in Fig. 3.
Figure 3: Unsolvable data construction workflow, where Fig. 8 demonstrates with an example.
@@
First, given a solvable question $\boldsymbol{x}_{i}$ along with the ground truth $\boldsymbol{y}_{i}$ from $\mathcal{D}_{a}$, we employ an instruction-tuned model $\mathcal{M}_{I}$ (GPT-4o (openai2024gpt4o, )) to extract 1–3 mathematical necessary conditions $\{{\boldsymbol{c}_{i}^{j}}\}_{j=1}=\mathcal{M}_{I}(\boldsymbol{x}_{i})$, with the number of extracted conditions varying according to actual conditions in $\boldsymbol{x}_{i}$. For each ${\boldsymbol{c}_{i}^{j}}$, we employ an advanced reasoning model $\mathcal{M}_{R}$ (Deepseek-R1 (deepseekai2025deepseekr1incentivizingreasoningcapability, )) to rewrite $\boldsymbol{x}_{i}$  [^1] We employ $\mathcal{M}_{I}$ to perform relatively simple tasks such as condition extraction, while $\mathcal{M}_{R}$ is utilized for complex analytical tasks with chain-of-thought (CoT) (NEURIPS2022_9d560961, ), ensuring both construction quality and efficiency., adhering to the two requirements in the rewriting instruction: (1) Remove the condition ${\boldsymbol{c}_{i}^{j}}$ from $\boldsymbol{x}_{i}$ or add a condition contradictory to ${\boldsymbol{c}_{i}^{j}}$ into $\boldsymbol{x}_{i}$, while keeping others unchanged, and (2) Ensure the rewritten problem is genuinely unsolvable. Two rewritten questions by removal and contradiction are separately generated for each ${\boldsymbol{c}_{i}^{j}}$ and several rewritten questions $\{{\boldsymbol{\hat{x}}_{i}^{j}}\}_{j=1}=\mathcal{M}_{R}(\boldsymbol{x}_{i},\boldsymbol{y}_{i})$ to the question $\boldsymbol{x}_{i}$ are obtained.
@@
Then, given $\boldsymbol{x}_{i}$ and $\boldsymbol{y}_{i}$, for each ${\boldsymbol{\hat{x}}_{i}^{j}}$, $\mathcal{M}_{I}$ is employed to compare ${\boldsymbol{\hat{x}}_{i}^{j}}$ and ${\boldsymbol{\hat{x}}_{i}}$ to verify whether ${\boldsymbol{\hat{x}}_{i}^{j}}$ only rewrites one condition in $\boldsymbol{x}_{i}$ to avoid other conditions modified, and $\mathcal{M}_{R}$ is utilized to analyze whether ${\boldsymbol{\hat{x}}_{i}^{j}}$ is indeed unsolvable. If both criteria are satisfied, ${\boldsymbol{\hat{x}}_{i}^{j}}$ is saved; otherwise, filtered out. For the saved ${\boldsymbol{\hat{x}}_{i}^{j}}$, both $\mathcal{M}_{I}$ and $\mathcal{M}_{R}$ are also required to output the rewritten condition ${\boldsymbol{\hat{c}}_{i}^{j}}=\mathcal{M}_{I}({\boldsymbol{\hat{x}}_{i}^{j}},\boldsymbol{x}_{i})$ and the unsolvable reason analysis ${\boldsymbol{\hat{a}}_{i}^{j}}=\mathcal{M}_{R}({\boldsymbol{\hat{x}}_{i}^{j}},\boldsymbol{x}_{i},\boldsymbol{y}_{i})$ to assist human evaluation on some challenging mathematical problems.
Table 2: Data statistics of $\mathcal{D}_{a}$, and the removal (Rm.) and contradiction (Cont.) questions after Step 1&2 and 3. In Step 3, we also present the number of problems in two difficulty levels ($d$=0 or $d$=1).
@@
$d$=0 $d$=1 Sum $d$=0 $d$=1 Sum $d$=0 $d$=1 Sum
AIME (aimo2024aime, ) 30 70 71 141 24 43 67 22 43 65 46 86 132
AMC (aimo2024amc, ) 83 185 192 377 47 84 131 60 104 164 107 188 295
MATH (hendrycks2021measuring, ) 100 233 216 449 77 77 154 75 89 164 152 166 318
Minerva (minervamath2024, ) 100 207 201 408 87 98 185 76 96 172 163 192 357
@@
Finally, we format $\{\boldsymbol{x}_{i},\boldsymbol{y}_{i},{\boldsymbol{\hat{x}}_{i}^{j}},{\boldsymbol{\hat{c}}_{i}^{j}},{\boldsymbol{\hat{a}}_{i}^{j}}\}$ in one sample and deliver to employed evaluators for assessment. Evaluators are required to determine whether ${\boldsymbol{\hat{x}}_{i}}^{j}$ meets the two criteria of model verification in Step 2. If verified, evaluators then need to rate the difficulty level ${{d}_{i}^{j}}$ of identifying the unsolvability to ${\boldsymbol{\hat{x}}_{i}^{j}}$: problems with obvious missing or contradictory conditions are regarded as simple cases to identify unsolvability and set ${{d}_{i}^{j}}=0$, while those requiring complex reasoning and analysis to determine are regarded as hard and set ${{d}_{i}}^{j}=1$. The validated and annotated question ${\boldsymbol{\hat{x}}_{i}^{j}}$ is then formatted and incorporated into the unsolvable dataset $\mathcal{D}_{u}$.
Both the solvable dataset $\mathcal{D}_{a}$ and the unsolvable dataset $\mathcal{D}_{a}$ are combined to constitute the ReliableMath dataset $\mathcal{D}_{r}$. Data statistics are in Table 2. More details of data construction and human check guidelines can be found in Appendix F and G.1 respectively.
@@ 4 Experiments
Table 3: Reliability evaluations of Precision (Prec.) / Prudence (Prud.) and lengths of generations (Len.) on Solvable ($\mathcal{A}$) and Unsolvable ($\mathcal{U}$) subsets on a series of reasoning and instruction LLMs employing the standard and reliable prompts respectively.
LLMs Prompt Solvable ($\mathcal{A}$) Unsolvable ($\mathcal{U}$)   Prec.   Prud.
Prec.($\mathcal{A}$) Prud.($\mathcal{A}$) Len. Prec.($\mathcal{U}$) Prud.($\mathcal{U}$) Len.
@@
DeepSeek-R1 standard 0.748 0.000 4.11k 0.000 0.000 6.50k 0.374 0.000
reliable 0.735 0.000 3.81k 0.549 0.007 4.40k 0.642 0.004
o3-mini standard 0.665 0.000 1.51k 0.001 0.000 4.90k 0.333 0.000
reliable 0.716 0.006 1.57k 0.293 0.005 4.20k 0.504 0.006
Distill-32B standard 0.712 0.000 4.99k 0.000 0.000 14.50k 0.356 0.000
reliable 0.684 0.000 5.05k 0.418 0.002 9.40k 0.551 0.001
Distill-14B standard 0.671 0.000 6.51k 0.000 0.000 16.00k 0.335 0.000
reliable 0.629 0.000 6.23k 0.465 0.001 11.00k 0.547 0.000
Distill-7B standard 0.610 0.000 6.16k 0.000 0.000 6.50k 0.305 0.000
reliable 0.575 0.000 6.24k 0.003 0.000 6.60k 0.289 0.000
Distill-1.5B standard 0.419 0.000 9.00k 0.000 0.000 9.60k 0.209 0.000
reliable 0.396 0.000 9.37k 0.000 0.000 9.70k 0.198 0.000
@@
DeepSeek-V3 standard 0.655 0.000 1.43k 0.001 0.000 1.90k 0.328 0.000
reliable 0.665 0.000 1.34k 0.377 0.003 1.50k 0.521 0.001
GPT-4o standard 0.460 0.000 0.72k 0.001 0.000 0.80k 0.230 0.000
reliable 0.460 0.006 0.58k 0.335 0.025 0.60k 0.397 0.015
Qwen2.5-7B standard 0.498 0.000 0.83k 0.000 0.000 0.90k 0.249 0.000
reliable 0.505 0.000 0.82k 0.027 0.000 0.90k 0.266 0.000
Qwen2.5-1.5B standard 0.450 0.000 0.71k 0.000 0.000 0.80k 0.225 0.000
reliable 0.422 0.000 0.74k 0.015 0.000 0.80k 0.218 0.000
@@
Experiments are conducted on a series of reasoning and instruction LLMs on ReliableMath dataset. Reasoning LLMs include DeepSeek-R1, DeepSeek-R1-Distill-Qwen (Distill) 32B, 14B, 7B, 1.5B (deepseekai2025deepseekr1incentivizingreasoningcapability, ), and o3-mini (openai2025o3mini, ). Instruction LLMs contain DeepSeek-V3 (deepseekai2024deepseekv3technicalreport, ), Qwen2.5-Math-Instruct (Qwen2.5) 7B, 1.5B (qwen2.5, ), and GPT-4o (openai2024gpt4o, ). We employ a standard mathematical problem-solving prompt and a reliable prompt that also allows identifying unsolvability or refusal. All generations are produced by greedy decoding. Model and prompt details are in Appendix E.2, H.1, and I.1.
4.1 Findings
@@
Experiments of LLMs’ reliability on mathematical reasoning tasks are presented in Table 3. Results on respective subsets are presented in Appendix H.2. Several findings are listed below.
@@
a. LLMs fail to directly identify the unsolvability of problems or refuse to answer but attempt to reason with substantial tokens, diminishing the reliability and aggravating the overthinking issue (wang2025harnessingreasoningeconomysurvey, ; chen2025think23overthinkingo1like, ; fan2025missingpremiseexacerbatesoverthinking, ). As in Table 3, all LLMs using standard prompts rarely exhibit the capability of identifying unsolvability or refusing, with both Prec.($\mathcal{U}$) and Prud.($\mathcal{U}$) close to 0. As in Fig. 4 (a), the generation lengths of unsolvable problems using standard prompts (usv std) are especially longer than others on four typically used LLMs. Upon examining the outputs, we found that while LLMs may recognize potential issues of unsolvable problems during reasoning, they will not explicitly acknowledge them. Instead, LLMs attempt to backtrack and reflect in reasoning, leading to excessively long results, or fabricate conditions to provide a hallucinated response, as demonstrated in Fig. 11. This behavior stems from the fact that during training, LLMs are encouraged to provide a certain answer and lack the ability to directly identify the unsolvability or refuse to answer.
Figure 4: Illustrations of (a) Lengths of generations on four representative LLMs, (b) Reliability figure of Precision on AIME, and (c) Reliability figure of Precision on MATH. In figure (a), slv and usl denote solvable and unsolvable problems, while std and real denote using standard and reliable prompts, respectively. Completed figures on all datasets are presented in Fig. 12. We employ different colors to indicate different LLMs. ◆ and ★ in figures represent using standard and reliable prompts. The dashed lines connecting two points in the same color represent the trajectory of reliability changes from the standard prompt ◆ to the reliable prompt ★ of an LLM.
@@
b. After employing reliable prompts, the reliability on solvable problems remains stable, but improves significantly on unsolvable problems, albeit still falling short of solvable problems. As in Table 3, after using reliable prompts, Precision on solvable problems remains and fluctuates within ±$\pm 0.08$ of Prec.($\mathcal{A}$), but significantly increases on unsolvable problems in approximately $0.2\sim 0.5$ of Prec.($\mathcal{U}$) on typically used LLMs like DeepSeek-R1 and DeepSeek-V3, while Prud.($\mathcal{A}$) and Prud.($\mathcal{U}$) also marginally improve. The observation implies that reliable prompts can alleviate the reliability issue on unsolvable cases without sacrificing the performance on solvable data. Moreover, in Fig. 4 (a), the sequence lengths after using reliable prompts also decrease noticeably (real vs. std), suggesting that reliable prompts can also mitigate the overthinking issue.
@@
c. For larger-size LLMs, the reliability of reasoning LLMs generally outperforms instruction LLMs when using reliable prompts. Conversely, smaller reasoning LLMs demonstrate inferior reliability to instruction LLMs. As in Table 3, DeepSeek-R1 and o3-mini present greater reliability compared to DeepSeek-V3 and GPT-4o in Precision. However, Distill-7B and Distill-1.5B demonstrate comparable or slightly weaker reliability than their counterparts, Qwen2.5-7B and Qwen2.5-1.5B, particularly in unsolvable problems of Prec.($\mathcal{U}$). Such relatively small reasoning LLMs are more prone to excessive invalid overthinking, with averaged generation lengths of almost 10k, severely undermining LLM reliability. In addition, for LLMs from the same family, larger LLMs of the same family exhibit superior reliability than small LLMs (Distill-32B vs. Distill-14B; Distill-7B vs. Distill-1.5B; Qwen2.5-7B vs. Qwen2.5-1.5B) for both reasoning and instruction LLMs.
@@
d. Despite exhibiting great performance on solvable problems, relatively small LLMs can hardly identify the unsolvability and exhibit no improvements even when employing reliable prompts. Whether small-sized reasoning LLMs (Distill-7B and Distill-1.5B) or the instruction LLMs (Qwen2.5-7B and Qwen2.5-7B) achieve Prec.($\mathcal{A}$) of $0.4\sim 0.6$ on solvable problems. However, on unsolvable problems, Prec.($\mathcal{U}$) is almost always zero, even when using reliable prompts. This may be due to small LLMs overfitting on compulsorily solving mathematical problems, leading to a decrease in general instruction following ability and thus failing to identify the unsolvability.
e. LLMs demonstrate greater reliability on relatively simple mathematical tasks than challenging ones after using reliable prompts. As in Fig. 4 (b) and (c), the Prec.($\mathcal{A}$) is distributed within $0.1\sim 0.6$ for more challenging olympic-level AIME, while Prec.($\mathcal{U}$) falls in $0.0\sim 0.4$. In contrast, for relatively simple high school-level MATH, Prec.($\mathcal{A}$) and Prec.($\mathcal{U}$) are approximately $0.7\sim 0.9$ and $0.0\sim 0.6$, suggesting LLMs are prone to identify unsolvable problems on MATH dataset.
@@
Figure 5: Illustrations of (a) Prec.($\mathcal{U}$) of unsolvable problems constructed by two rewriting types in Sec. 3.2 (removal & contradiction), and (b) Prec.($\mathcal{U}$) of unsolvable problems of two difficulty levels rated by human experts in Sec. 3.3 (0: simple & 1: hard) on several typically used LLMs employing reliable prompts. Completed results on all LLMs are presented in Fig. 13 in Appendix.

### v2

@@ 3.3 Unsolvable Data Construction Workflow
To obtain high-quality unsolvable problems by the rewriting schemes, further professional verification and refinement are required. Prior work on creating unanswerable problems in knowledge-intensive tasks has relied exclusively on manual collection or filtering (amayuelas-etal-2024-knowledge; yin_large_2023). However, mathematical problems present greater challenges, and fully manual rewriting and validation are prohibitively expensive. Therefore, we design a three-stage data construction workflow to ensure both efficiency and quality, as in Fig. 3.
Figure 3: Unsolvable data construction workflow, where the example version is presented in Supplement 12.
@@
[…]
@@
Then, given $\boldsymbol{x}_{i}$ and $\boldsymbol{\hat{y}}_{i}$, for each $\boldsymbol{\tilde{x}}_{i}^{j}$}, $\mathcal{M}_{I}$ is employed to compare $\boldsymbol{\tilde{x}}_{i}^{j}$} and $\boldsymbol{x}_{i}$} to verify whether $\boldsymbol{\tilde{x}}_{i}^{j}$} only rewrites one condition in $\boldsymbol{x}_{i}$ to avoid other modified conditions, and $\mathcal{M}_{R}$ is utilized to analyze whether $\boldsymbol{\tilde{x}}_{i}^{j}$} is indeed unsolvable. If both criteria are satisfied, $\boldsymbol{\tilde{x}}_{i}^{j}$} is retained; otherwise, filtered out. For retained $\boldsymbol{\tilde{x}}_{i}^{j}$}, both $\mathcal{M}_{I}$ and $\mathcal{M}_{R}$ are also required to output the rewritten condition ${\boldsymbol{\tilde{c}}_{i}^{j}}=\mathcal{M}_{I}({\boldsymbol{\tilde{x}}_{i}^{j}},\boldsymbol{x}_{i})$ and the unsolvable reason analysis ${\boldsymbol{\tilde{a}}_{i}^{j}}=\mathcal{M}_{R}({\boldsymbol{\tilde{x}}_{i}^{j}},\boldsymbol{x}_{i},\boldsymbol{\hat{y}}_{i})$ to assist human evaluation in the next stage.
@@
dd=0 dd=1 Sum dd=0 dd=1 Sum dd=0 dd=1 Sum
AIME 30 70 71 141 24 43 67 22 43 65 46 86 132
AMC 83 185 192 377 47 84 131 60 104 164 107 188 295
MATH 100 233 216 449 77 77 154 75 89 164 152 166 318
Minerva 100 207 201 408 87 98 185 76 96 172 163 192 357
@@
Table 2: Data statistics of $\mathcal{D}_{a}$, and the removal (Rm.) and contradiction (Cont.) questions after Step 1&2 and 3. In Step 3, we also present the number of problems in two difficulty levels (dd=0 or dd=1).
@@
Finally, we format all $\{\boldsymbol{x}_{i},\boldsymbol{y}_{i},{\boldsymbol{\tilde{x}}_{i}^{j}},{\boldsymbol{\tilde{c}}_{i}^{j}},{\boldsymbol{\tilde{a}}_{i}^{j}}\}$ and deliver them to employed experts holding a master’s degree or higher in STEM field for assessment. Experts first determine whether ${\boldsymbol{\tilde{x}}_{i}}^{j}$ meets the two criteria of model verification in Step 2. If verified, they then rate the difficulty level ${d}_{i}^{j}$} of identifying the unsolvability to $\boldsymbol{\hat{x}}_{i}^{j}$}: problems with obvious missing or contradictory conditions are regarded as simple cases to identify unsolvability and set ${{d}_{i}^{j}}=0$, while those requiring step-by-step reasoning and analysis to determine are regarded as hard with ${{d}_{i}}^{j}=1$. The validated question $\boldsymbol{\hat{x}}_{i}^{j}$} is then formatted and incorporated into the unsolvable dataset $\mathcal{D}_{u}$.
Both $\mathcal{D}_{a}$ and $\mathcal{D}_{a}$ are combined to constitute the ReliableMath dataset $\mathcal{D}_{r}$. Data statistics are presented in Table 2. Definitions of notations are detailed in Supplement 8. More details of data construction and human check guidelines can be found in Supplement 10-12. Specifically, we specify detailed principles for difficulty level annotation to minimize variations arising from the subjectivity of human experts in Supplement 12.3.
@@ 4 Experiments
LLMs Prompt Solvable ($\mathcal{A}$) Unsolvable ($\mathcal{U}$) Succ. Refu.
Succ.($\mathcal{A}$) Refu.($\mathcal{A}$) Len. Succ.($\mathcal{U}$) Refu.($\mathcal{U}$) Len.
$s_{p}$ $s_{o}$ $s_{p}$ $s_{o}$
@@
DeepSeek-R1 standard 73.81 74.76 0.00 4.09k 59.80 0.00 0.00 6.51k 52.08 0.00
reliable 73.17 73.49 0.00 3.82k 70.79 54.89 1.76 4.38k 68.08 0.88
o3-mini standard 64.85 66.44 0.00 1.49k 14.61 0.08 0.00 5.03k 36.50 0.00
reliable 69.95 71.58 0.64 1.56k 25.78 29.31 0.87 4.18k 49.15 0.75
Distill-32B standard 69.32 71.24 0.00 4.99k 31.68 0.00 0.00 14.52k 43.06 0.00
reliable 67.40 68.36 0.00 5.05k 51.53 41.87 0.24 9.36k 57.29 0.12
Distill-14B standard 66.12 67.08 0.00 6.51k 36.94 0.00 0.00 16.97k 42.52 0.00
reliable 62.30 62.93 0.00 6.24k 59.61 46.45 0.12 10.94k 57.83 0.06
Distill-7B standard 60.37 61.02 0.00 6.16k 1.74 0.00 0.00 6.50k 30.77 0.00
reliable 57.20 57.52 0.00 6.22k 1.99 0.26 0.00 6.59k 29.24 0.00
Distill-1.5B standard 40.26 41.86 0.00 9.02k 2.74 0.00 0.00 9.55k 21.20 0.00
reliable 38.98 39.62 0.00 9.35k 2.18 0.00 0.00 9.69k 20.21 0.00
@@
DeepSeek-V3 standard 64.22 65.49 0.00 1.43k 27.68 0.09 0.00 1.90k 39.37 0.00
reliable 65.50 66.45 0.00 1.35k 43.29 37.75 0.86 1.54k 53.25 0.43
GPT-4o standard 42.82 46.01 0.00 0.72k 10.24 0.10 0.00 0.78k 24.79 0.00
reliable 42.16 46.01 0.64 0.57k 27.34 33.50 6.95 0.65k 37.23 3.79
Qwen2.5-7B standard 44.73 49.84 0.00 0.86k 1.43 0.00 0.00 0.87k 24.00 0.00
reliable 44.73 50.47 0.00 0.83k 3.38 2.73 0.00 0.88k 25.32 0.00
Qwen2.5-1.5B standard 41.22 45.05 0.00 0.71k 1.89 0.00 0.00 0.72k 22.04 0.00
reliable 39.63 42.18 0.00 0.72k 2.17 1.44 0.00 0.75k 21.36 0.00
Table 3: Reliability performance of Success Rate (Succ.) / Refusal Rate (Refu.) and Response Length (Len.) on Solvable ($\mathcal{A}$) and Unsolvable ($\mathcal{U}$) subsets on a series of reasoning and instruction LLMs.
4.1 Evaluation Settings
Models
@@ 4.1 Evaluation Settings
Experiments are conducted on a series of reasoning and instruction LLMs on ReliableMath dataset. Reasoning LLMs include DeepSeek-R1, R1-Distill-Qwen (Distill) 32B, 14B, 7B, 1.5B (deepseekai2025deepseekr1incentivizingreasoningcapability), and o3-mini (openai2025o3mini). Instruction LLMs contain DeepSeek-V3 (deepseekai2024deepseekv3technicalreport), Qwen2.5-Math-Instruct (Qwen2.5) 7B, 1.5B (qwen2.5), and GPT-4o (openai2024gpt4o). We employ standard math problem-solving prompts and reliable prompts that also allows identifying unsolvability or refusal. All generations are produced by greedy decoding. Model and prompt details are in Supplement 13 and 15.
Evaluation Metrics
Following Sec. 2, we employ two metrics to evaluate LLM reliability on mathematical reasoning tasks: Success Rate (Succ.) and Refusal Rate (Refu.). Succ. measures the proportion of successful responses $\mathcal{S}$ calculated on solvable ($\mathcal{A}$) and unsolvable ($\mathcal{U}$) questions separately, and then averaging as:
$\vskip-22.76219pt\textit{Succ.}(\mathcal{A})$ $=\frac{\#\mathcal{AS}}{\#\mathcal{A}},\ \ \textit{Succ.}(\mathcal{U})=\frac{\#\mathcal{US}}{\#\mathcal{U}}$ (1)
Succ. $=\frac{1}{2}\left[\textit{Succ.}(\mathcal{A})+\textit{Succ.}(\mathcal{U})\right].$ (2)
Refu. represents the proportion of refused responses $\mathcal{R}$ on $\mathcal{A}$ and $\mathcal{U}$ as follows.
$\vskip-22.76219pt\textit{Refu.}(\mathcal{A})$ $=\frac{\#\mathcal{AR}}{\#\mathcal{A}},\ \ \textit{Refu.}(\mathcal{U})=\frac{\#\mathcal{UR}}{\#\mathcal{U}}$ (3)
Refu. $=\frac{1}{2}\left[\textit{Refu.}(\mathcal{A})+\text{{Refu.}}(\mathcal{U})\right]$ (4)
When assessing LLMs’ reliability, Succ. is prioritized, and the higher the Succ., the better the reliability. When achieving comparable Succ. for two LLMs, Refu. is then considered, with a higher Refu. being preferable, as in Def. 1.
[…]
4.2 Experimental Findings
@@
Experiments of LLMs’ reliability on mathematical reasoning tasks are presented in Table 3. Completed results on respective subsets are in Supplement 13. Several findings are listed below.
@@
a. LLMs can occasionally recognize the unreasonableness or illogicality of unsolvable problems, but fail to directly identify the unsolvability or refuse to answer but attempt to fabricate reasoning steps with substantial tokens, diminishing the reliability and aggravating the overthinking issue (wang2025harnessingreasoningeconomysurvey; chen2025think23overthinkingo1like; fan2025missingpremiseexacerbatesoverthinking). As in Table 3, all LLMs using standard prompts rarely exhibit the capability of explicitly identifying unsolvability or refusing, with both $s_{o}$ of Succ.($\mathcal{U}$) and Refu.($\mathcal{U}$) close to 0. The generation lengths of unsolvable problems using standard prompts are especially lengthy. In Table 4, after examining the outputs, we list the keywords related to deep thinking, backtrack or reflection behaviors, and typical patterns to recognize unsolvability of “removal” and “contradiction” problems respectively, suggesting that LLMs may recognize potential issues of unsolvable problems during reasoning but still fabricate reasonings to provide a hallucinated response. This behavior stems from the fact that during training, LLMs are always encouraged to provide a certain answer given any questions. Hence they lack the ability to critically identify the unsolvability, making their reliability are highly susceptible to attack.
All Remove Contradict
wait 59.59 lack 4.00 contradict- 4.44
alternative 13.60 los- 3.85 opposite 4.25
correct 8.12 mistake 3.05 mistake 3.73
again 4.86 assum- 2.87 incorrect 1.92
change 4.73 miss- 1.83 inconsisten- 0.83
check 4.58 undefine- 0.97 conflict 0.76
Table 4: The keywords statistics and their frequencies in the responses on all models and datasets using standard prompts, where “All” represents the keywords related to thinking or reflection behavior on all unsolvable problems, while “Remove” and “Contradict” denote the respective keyword patterns to identify unreasonable problems in the two types of unsolvable problems. pre- denotes the words sharing the prefix pre*. For example, assum- includes assume, assumption, and assuming.
Figure 4: Results of Success Rate (Succ.) on o3-mini on different test sets (AIME, AMC, MATH, Minerva) on both solvable (slv) and unsolvable (usl) subsets using standard (std) and reliable (real) prompts, respectively.
@@
b. After employing reliable prompts which enable LLMs to critically identify the solvability or refuse to answer, the reliability on solvable problems remains stable, but improves significantly on unsolvable problems, albeit still falling short of solvable problems. As in Table 3, after using reliable prompts, Succ.($\mathcal{A}$) on solvable problems remains or slightly fluctuates due to prompt sensitivity, but significantly increases on unsolvable problems in approximately $20\sim 50$ in $s_{p}$ and $s_{o}$ of Succ.($\mathcal{U}$) on typically used LLMs like DeepSeek-R1 and GPT-4o, while Refu.($\mathcal{A}$) and Refu.($\mathcal{U}$) also marginally improve. The observation implies that reliable prompts can alleviate the unreliability issue on unsolvable cases without sacrificing the performance on solvable data. Moreover, the sequence lengths after using reliable prompts also decrease noticeably, suggesting that reliable prompts can also mitigate the overthinking issue.
@@
c. For larger-size LLMs, the reliability of reasoning LLMs generally outperforms instruction LLMs when using reliable prompts. Conversely, smaller reasoning LLMs demonstrate inferior reliability to instruction LLMs. As in Table 3, DeepSeek-R1 and o3-mini present greater reliability compared to DeepSeek-V3 and GPT-4o in Succ.($\mathcal{U}$). However, Distill 7B and 1.5B demonstrate slightly weaker reliability than their counterparts, Qwen2.5 7B and 1.5B, particularly in Refu.($\mathcal{U}$). Such relatively small reasoning LLMs are prone to excessive invalid overthinking, with averaged generation lengths over 10k, severely undermining LLM reliability. In addition, for LLMs from the same family, larger LLMs exhibit superior reliability than small LLMs (Distill 32B vs. 14B; Distill 7B vs. 1.5B; Qwen2.5 7B vs. 1.5B) for both reasoning and instruction LLMs.
@@
d. Despite performing well on solvable problems, relatively small LLMs can hardly detect unsolvability and show no improvement even when using reliable prompts. Whether small-sized reasoning LLMs (Distill 7B and 1.5B) or the instruction LLMs (Qwen2.5 7B and 1.5B) achieve Succ.($\mathcal{A}$) of $40\sim 60$ on solvable problems, but Succ.($\mathcal{U}$) are always almost 0 on unsolvable problems even using reliable prompts. This is because small LLMs after training overfit to generate certain answers to any mathematical problems even for unsolvable ones, thereby weakening their general instruction-following ability and preventing them from recognizing unsolvable cases.
e. For mathematical problems at different levels, LLMs exhibit high sensitivity on solvable problems but remain invariance on unsolvable ones. As in Fig. 4, for solvable problems (slv), LLMs achieve higher Succ.($\mathcal{A}$) on relatively simple high school–level MATH problems compared with the more challenging Olympiad-level AIME and AMC, and the college-level Minerva, employing either standard or reliable prompts. In contrast, for unsolvable problems (usl), Succ.($\mathcal{U}$) remain comparable across different levels, suggesting that the ability of LLMs to identify an unsolvable problem is independent of the difficulty level of the corresponding original solvable problem.
@@
Figure 5: Illustrations of Succ.($\mathcal{U}$) of unsolvable problems regarding (a) two rewriting schemes (removal & contradiction), and (b) two difficulty levels labeled by experts (0: simple & 1: hard) on several typically used LLMs using reliable prompts. Completed results on all LLMs are presented in Supplement C.3.